find three consecutive integers such that the sum of the first and twice the second is 17 more than twice the third.
Answer by Paul(988)

Let the three consecutive numbers be, x, x+1, and x+2
Twice the second = 2(x+1)
"is" =
17 more than twice the third = 17+2(x+2)
Add them togeher.
x+2(x+1)=17+2(x+2)
x+2x+2=17+2x+4
3x+2=2x+21 --->SImplfy
x=19
19+1=20
19+2=21

Hence the three numbers are 19, 20, and 21.
Paul.